Simone,Alfio,JuanCArlos, Lorenzo

OPTICAL ALIGNMENT OF W2

Checking the fault of the last day. This fault was related to the continous shut of the ccd39 BCU. We suspect that the problem could be in the power connector.

To acces the power connector of the device, the "black box" has to be take off the W Unit. Looking the connector and investiating the fault, we can see that the connector has a L clamp, this clamp seemsthat isnot doing anything, after moving the connector many times the BCU didn't shut all the times, so wechecked the relay connector.
The relay connector is bypassed and we have seen that the BCU didn't shut for many minutes. Then we decided tomount the system in the black box again and mount in the W Unit.

The mount in the W unit, we crossed with a problem ofone tube for the cooling plate, the connector couldn't be connected and it looked brocken. We make a new connector with a bit longer tube and we could connect it ( one the reason why it was too strong to plug it, it was that the cool plate keep some pressure inside, so to plut again is necessary to forceit a bit, this was done in a safety conditions with the cool plate on the table and a easy acces to the plugs).

After to put the "black box" in the W unit, we started to solve a bug with the measurements of the pupil diameters and lato. Looking the software code we found the bug and the measures of the pupils are more realistics now.

It still necesary toimprove the alignent to reach the diameter of the pupil 31 pixels and the latto at 36.

-- JuanCarlosGuerra - 2009-09-03
Topic revision: r2 - 03 Sep 2009, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback